OpenStudy (anonymous):

x + p is a factor of both; ax^2 + b and ax^2 + bx − ac. (i) Show that p^2=-b/a and that p=-b-ac/b (ii) Hence show that p^2 + p^3 = c

OpenStudy (anonymous):

(I)- (x+p) is factor, therefor f(x) = 0 when x=-p(factor theorem). So just sub in -p into each equation and rearrange till you get required result. A(-p)^2+b=> p^2=-b/a A(-p)^2 +b(-p) -ac => p=(-b-ac)/b (ii) p^2 = -b/a, p=(-b-ac)/b P^3= p x p^2 => (-b-ac)/b + [(-b/a) x (-b-ac)/b] (this is p^2 + p^3) Multiply out and simplify so it equals C